Consider the following information for Watson Power Co.: Debt: 4,500 8 percent coupon bonds outstanding, $1,000 par value, 18 years to maturity, selling for 104 percent of par; the bonds make semiannual payments. Common stock: 112,500 shares outstanding, selling for $63 per share; the beta is 1.15. Preferred stock: 13,500 shares of 7 percent preferred stock outstanding, currently selling for $106 per share. Market: 10 percent market risk premium and 6 percent risk-free rate. Assume the company's tax rate is 35 percent. Find the WACC.

Multiple Choice

  • 12.03%

  • 11.86%

  • 12.26%

  • 12.8%

  • 11.76%
Answer #1

rate positively ..

Rate of equity =
Using CAPM rate of equity = Risk free rate + market risk premium * beta
=6%+10%*1.15
17.500%
Rate of debt (after tax)
we have to use financial calculator to compute YTM
Put in calcualtor
PV -1040
FV 1000
PMT 1000*8%/2 40
N 18*2 36
Compute I 3.79%
YTM = 7.59%
tax rate = 35%
therefore rate of debt (after tax) = 4.93%
rate of debt (after tax) = 4.93%
rate of preferred stock = Annual dividend/Current price
=7/106
6.60%
Market value weight Cost of capital weight * cost
Source
equity 7087500 =112500*63 53.70% 17.50% 9.40%
debt 4680000 =4500*1000*104% 35.46% 4.93% 1.75%
preferred stock 1431000 =13500*106 10.84% 6.60% 0.72%
13198500 11.86%
Ans = 11.86%